OpenStudy (anonymous):

x^3 – 2^3 = (x – 2)(x^2 + 2x + 2^2) can someone explain this?

hero (hero):

\[(x - 2)(x^2 + 2x + 4) \\= x^2(x - 2) + 2x(x - 2) + 4(x - 2) \\= x^3 - 2x^2 + 2x^2 - 4x + 4x - 8 \\ = x^3 - 8\]

hero (hero):

@dabeezy